ow many different ID cards can be made if there are 6 digits on a card and no digit can be used more than once?
1
Expert's answer
2020-12-17T18:53:36-0500

Permutation "_{10}P_6=\dfrac{10!}{(10-6)!}=10(9)(8)(7))6)(5)=151200"